OpenStudy (anonymous):

Factorise x^2 - 2X -15

OpenStudy (lgbasallote):

hint break down the middle term x^2 - 5x + 3x - 15

OpenStudy (lgbasallote):

does that help?

OpenStudy (anonymous):

how does that help? Are you suppose to complete the square?

OpenStudy (anonymous):

That does'nt help at all!

mathslover (mathslover):

\[\huge{x^2-2x-15}\] \[\huge{x^2+3x-5x-15}\] \[\huge{x(x+3)-5(x+3)}\] can u do further ?

mathslover (mathslover):

Note that : \[\large{a(b+c)-d(b+c)=(a-d)(b+c)}\] This property is known as : inverse of distributive property

OpenStudy (cherylim23):

Alternatively, complete the square. It works all the time, despite getting fractions at some other times. x^2 - 2x -15 = (x- 2/2)^2 - (2/2)^2 - 15 = (x-1)^2 - 1 - 15 = (x-1)^2 - 16 Use difference of two squares, = (x-1+4)(x-1-4) = (x+3)(x-5)

OpenStudy (anonymous):

x^2 - 2X -15 (x+a)(x+b) a and b are two numbers that when you multiply them together you get -15 and when you add them together you -2

OpenStudy (anonymous):

Simple as that